Opened 10 years ago

Last modified 9 years ago

#477 new flight processing

RG12/09, flight day 120b/2013, Wessex

Reported by: knpa Owned by: anhi
Priority: alpha 4 high Milestone:
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by besm)

Data location: /users/rsg/arsf/arsf_data/2013/flight_data/uk/RG12_09-2013_120_Wessex/

Data arrived from ARSF via network transfer on 13/05/2013

RG12/09 Scientific objective: Biodiversity and the provision of multiple ecosystem services in current and future lowland multifunctional landscapes

RG12/09 Priority: 8+

PI: France Gerard

Sensors:

Camera (Delivered 17/07/2013)
Eagle (Delivered 17/07/2013)
Hawk (Delivered 17/07/2013)
LiDAR (Delivered 17/07/2013)

Change History (29)

comment:1 Changed 10 years ago by knpa

  • Summary changed from RG12/09, flight day 120/2013, Wessex to RG12/09, flight day 120b/2013, Wessex

comment:2 Changed 10 years ago by anhi

Finished navigation processing

comment:3 Changed 10 years ago by knpa

  • Description modified (diff)

comment:4 Changed 10 years ago by knpa

  • Description modified (diff)

comment:5 Changed 10 years ago by knpa

  • Priority changed from alpha 4 medium to alpha 4 high

comment:6 Changed 10 years ago by anhi

Hawk header files contain errors so copies have been made and the start and stop times and positions have been corrected to match the eagle files.

comment:7 Changed 10 years ago by anhi

SCT Values

Eagle Hawk
1 1.85 1.46
2 1.73 -0.58
3 2.88 0.51
4 2.48 0.08

comment:8 Changed 10 years ago by anhi

Hyperspectral delivery created

comment:9 Changed 10 years ago by anhi

LiDAR delivery created:

Pitch:-0.003136
Roll:-0.00121

comment:10 Changed 10 years ago by tipo

LiDAR delivery checked, no issues found.

comment:11 Changed 10 years ago by anhi

RCD Delivery Created.

comment:12 Changed 10 years ago by anhi

Hyperspectral delivery created with new hawk files.

comment:13 Changed 10 years ago by knpa

Beginning HS delivery check.

comment:14 Changed 10 years ago by knpa

Delivery check complete. All looks good, however spotted underflow which you should mention on the Hawk level 1 files - mostly on the latter bands, towards the top of the line.

Don't forget to zip up the mapped files.

comment:15 Changed 10 years ago by besm

DC for RCD complete - we concluded we should mention the missing images in the readme. Otherwise looks fine.

comment:16 Changed 10 years ago by anhi

Mentioned the image coverage in the Readme.

comment:17 Changed 10 years ago by besm

DC for RCD complete. Looks fine.

comment:18 Changed 10 years ago by anhi

HS delivery with vectors and underflow mentioned created.

comment:19 Changed 10 years ago by besm

Delivery checked hyperspectral data. Now looks good, ready to deliver.

comment:20 Changed 10 years ago by besm

Updated LiDAR readme, need someone to delivery check.

comment:21 follow-up: Changed 10 years ago by tipo

Checked readme. No problems found.

comment:22 in reply to: ↑ 21 Changed 10 years ago by besm

Replying to tipo:

Checked readme. No problems found.

(Clarification: he means the LiDAR readme.)

comment:23 Changed 10 years ago by besm

Moving delivery data for camera, eagle, hawk, and LiDAR onto usb disk now.

comment:24 Changed 10 years ago by besm

Delivered hyperspectral, LiDAR and camera data on disk 161 to France Gerard at the Centre of Ecology Hydrology at OX108BB. 17 July 2013.

comment:25 Changed 10 years ago by besm

  • Description modified (diff)

comment:26 Changed 9 years ago by besm

Resent hyperspectral, LiDAR and camera data on disk 168 to France Gerard at same address. 08 August 2013.

comment:27 Changed 9 years ago by emca

  • Component changed from Processing: general to Archiving

comment:28 Changed 9 years ago by emca

Data ready for archiving. Email sent to Wendy@NEODC 01/10/2013.

comment:29 Changed 9 years ago by tipo

Split each hyperspectral and lidar line into 3 sections and redelivered.

Note: See TracTickets for help on using tickets.